# Variable Inertia Model Predictive Control for Fast Bipedal Maneuvers

### v1

@@ Abstract
This paper proposes a novel control framework for agile and robust bipedal locomotion, addressing model discrepancies between full-body and reduced-order models. Specifically, assumptions such as constant centroidal inertia have introduced significant challenges and limitations in locomotion tasks. To enhance the agility and versatility of full-body humanoid robots, we formalize a Model Predictive Control (MPC) problem that accounts for the variable centroidal inertia of humanoid robots within a convex optimization framework, ensuring computational efficiency for real-time operations. In this formulation, we incorporate a centroidal inertia network designed to predict the variable centroidal inertia over the MPC horizon, taking into account the swing foot trajectories—an aspect often overlooked in ROM-based MPC frameworks. Moreover, we enhance the performance and stability of locomotion behaviors by synergizing the MPC-based approach with whole-body control (WBC). The effectiveness of our proposed framework is validated through simulations using our full-body humanoid robot, DRACO 3 [1], demonstrating dynamic behaviors.
@@ I INTRODUCTION
The capability to perform highly dynamic (high-speed) locomotion for humanoid robots is one of the key elements towards achieving human-level task completion speed. However, controlling highly dynamic bipedal locomotion is challenging due to its nonlinear, underactuated, and hybrid nature. To handle these complexities, reduced-order models (ROMs) focusing on the centroidal dynamics [2] have been widely adopted, one prominent example being the single-rigid body (SRB) model [3].
Model predictive control (MPC) assuming the SRB model has been one of the dominant control strategies applied in quadrupedal locomotion [3, 4, 5]. More recently, this approach has been adapted in humanoid robots [6, 7, 8, 9, 10]. A commonality across these applications is their focus in optimizing locomotion variables such as reaction forces or footsteps under the assumption that the robot can be modeled as a single lumped mass — often informally referred to as the “potato model.” However, in stark contrast to most quadruped robots, this simplification falls short for humanoid robots that feature heavy legs. Specifically, significant changes in the rotational inertia tensor, e.g., resulting from leg swings, compromise the model’s accuracy. Since the MPC’s performance is highly dependent on the precision of the underlying model, the SRB model-based MPC controller becomes less suitable for humanoid robots undergoing high inertia variations.
@@
In this work, we present an MPC approach that considers the changes in centroidal rotational inertia due to different leg swing motions. Similar to our previous work [12], we construct a centroidal composite inertia neural network (CCINN) offline. This NN accurately captures the full-body composite rigid body inertia, eliminating the dependency on the joint configuration. Therefore, this seamlessly allows for the integration with the SRB model online. Unlike [12], which often faced challenges with poor local minima and slow computation time, we formulate our MPC problem to be convex for real-time applicability. Moreover, we incorporate the whole-body orientation (WBO) [13] dynamics of the robot into the SRB model. This approach excels the traditional use of the base link orientation [6, 7, 9], providing a more accurate dynamics representation of the SRB model. With these efforts to reduce the model discrepancy, our MPC predicts optimal reaction wrenches that serve as a more dynamically accurate reference to the low-level whole-body controller (WBC). This, in turn, enables more dynamic locomotion behaviors, including longer steps, higher walking speeds, and increased turning agility.
The organization of the remainder of this paper is as follows: In Section II, we revisit the dynamics models for control and formally define our problem. Section III details our approach to solving this problem through the MPC and WBC framework. Then, we present a set of numerical simulation results on the MPC in Section IV. Lastly, Section V concludes the paper with a short discussion and directions for future work.
@@ II PRELIMINARIES
II-A Full-body Dynamic Model
In this paper, we employ the DRACO $3$ humanoid robot [1], which is a $39~{}$\mathrm{k}\mathrm{g}$$ robotic system featuring rolling contact joints in its lower body, allowing for a broad spectrum of movements. With a total of $n_{a}=25$ degrees of freedom (DoF), DRACO $3$ comprises of $6$ DoF per leg, $6$ DoF per arm, and $1$ DoF in the neck. The full-body dynamics (FBD) of DRACO $3$ is given by:
$\mathbf{M}(\mathbf{q})\ddot{\mathbf{q}}+\mathbf{b}(\mathbf{q},\mathbf{\dot{q}})=\mathbf{S}_{a}^{\top}\bm{\tau}+\mathbf{J}_{\textrm{int}}(\mathbf{q})^{\top}\mathbf{F}_{\textrm{int}}+\mathbf{J}_{c}(\mathbf{q})^{\top}\mathbf{F}_{r},$ (1)
where $\mathbf{q}=[\mathbf{q}_{f}^{\top},\mathbf{q}_{a}^{\top},\mathbf{q}_{v}^{\top}]^{\top}\in\mathcal{Q}$, denotes the robot’s configuration, $\mathbf{q}_{f}\in SE(3)$ represents the position and orientation of the system’s floating base, $\mathbf{q}_{a}\in\mathbb{R}^{n}_{a}$ is the actuated joints’ configuration, $\mathbf{q}_{v}\in\mathbb{R}^{n}_{v}$ the unactuated joints’ configuration, and $\mathbf{\dot{q}}\in\mathbb{R}^{n_{a}+n_{v}+6}$ the system’s generalized velocity. $\mathbf{M}(\mathbf{q})\in\mathbb{S}_{\geq 0}^{n_{t}}$, and $\mathbf{b}(\mathbf{q},\dot{\mathbf{q}})\in\mathbb{R}^{n_{t}}$ denote the mass/inertia matrix and the sum of the centrifugal/Coriolis and gravitational force, respectively, where $n_{t}=n_{a}+n_{v}+6$. For simple notations, we use $\mathbf{M}$ and $\mathbf{b}$ instead of using $\mathbf{M}(\mathbf{q})$ and $\mathbf{b}(\mathbf{q},\dot{\mathbf{q}})$, respectively. The selection matrix for the actuated joint torque vector $\bm{\tau}\in\mathbb{R}^{n_{a}}$ is represented by $\mathbf{S}_{a}\in\mathbb{R}^{n_{a}\times n_{t}}$. In addition, $\mathbf{F}_{r}\in\mathbb{R}^{n_{c}}$ represents the concatenation of the $6$-dimensional contact wrenches, and $\mathbf{J}_{c}(\mathbf{q})\in\mathbb{R}^{n_{c}\times n_{t}}$ is the corresponding contact Jacobian. $\mathbf{F}_{\textrm{int}}\in\mathbb{R}^{n_{i}}$ and $\mathbf{J}_{\textrm{int}}(\mathbf{q})\in\mathbb{R}^{n_{i}\times n_{t}}$ represent the internal forces and their corresponding internal constraint Jacobian, respectively. For more details about the internal constraints, the reader is referred to [1].
@@ II-B Single Rigid Body Dynamic Model
ROMs offer the advantage of capturing the simplicity of locomotion principles while enabling the stabilization of robotic systems in fast and reactive manners. The SRB model has been among the most widely used to perform legged robot locomotion. The SRB model includes three major assumptions: 1) the leg dynamics have a negligible impact on the centroidal states of the robot, 2) the utilization of a constant rotational inertia matrix, $\mathbf{I}\in\mathbb{R}^{3\times 3}$, calculated based on the robot’s nominal configuration, and 3) the dynamics of the robot’s base orientation fully capture its overall orientation dynamics. Based on these assumptions, the equations of motion of the SRB model and its rotational motion are expressed by:
@@
where $m$ represents the lumped mass of the robot, and $\mathbf{p}_{c}\in\mathbb{R}^{3}$ denotes the position of the robot’s Center of Mass (CoM). $\mathbf{g}\in\mathbb{R}^{3}$ is the gravitational acceleration. $\mathbf{f}_{i}\in\mathbb{R}^{3}$ and $\bm{\tau}_{i}\in\mathbb{R}^{3}$ represent the $i^{\textrm{th}}$ ground reaction force and torque, respectively. $\mathbf{R}\in SO(3)$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the rotation matrix, which transforms from the body frame to world coordinates, and the body angular velocity. $\mathbf{r}_{i}\in\mathbb{R}^{3}$ is the $i^{\textrm{th}}$ footstep position relative to CoM. Note that all the quantities are expressed in the inertial frame unless explicitly stated otherwise.
@@ II-C Whole-body Orientation
where $\mathbf{I}_{G}$, and $\mathbf{k}_{G}$ represent the the centroidal composite rigid body inertia (CCRBI) [14] and centroidal angular momentum, respectively.
@@ III THE PROPOSED APPROACH
This section introduces our proposed variable inertia MPC-based control framework designed to enable robust dynamic locomotion of humanoid robots, as depicted in Fig 2.
@@ III-A Simplified Model
Inspired by the work of quadruped locomotion control [3, 15], we model the robot as an SRB subject to contact patches at each stance foot. However, there are two key differences in our approach: first, we explicitly take the full-body composite inertia into account along with the SRB. Additionally, instead of the robot’s torso orientation, we employ the WBO dynamics for the SRB orientation dynamics. Therefore, we rewrite the SRB rotational dynamics equations (3) and (4) in II-B as follows:
@@
where $\mathbf{I}(\mathbf{q})\in\mathbb{R}^{3\times 3}$ denotes the centroidal composite rigid body inertia (CCRBI) [2], which is a function of joint configuration $\mathbf{q}$. $\mathbf{R}_{\textrm{WBO}}$ represents the rotational centroidal orientation of a multibody system, while $\bm{\omega}_{\textrm{WBO}}\in\mathbb{R}^{3}$ denotes the system’s whole body angular velocity. For state-space representation, the rotational kinematics in (9) is rewritten using Euler angles as follows:
@@
where $\bm{\Theta}_{\textrm{WBO}}=[\theta,\phi,\psi]$ follows an earth-fixed roll($\theta$), pitch($\phi$), yaw($\psi$) convention, and $\mathbf{T}(\bm{\Theta}_{\textrm{WBO}})\in\mathbb{R}^{3\times 3}$ is the matrix that maps from Euler angle rates to angular velocities.
Assumption 1.
Assumption 2.
In our problem, we assume the matrix $\mathbf{T}$ is always invertible, namely $\phi\neq\frac{\pi}{2}$ (robot’s WBO is not pointed vertically, which in practice does not happen)
@@
We define the state variable $\bm{x}=[\bm{\Theta}_{\textrm{WBO}}^{\top},\bm{p}_{c}^{\top},\bm{w}_{\textrm{WBO}}^{\top},\dot{\bm{p}}_{c}^{\top}]^{\top}\in\mathbb{R}^{n_{x}}$ and the control input $\bm{u}=[\bm{\tau}_{1}^{\top},\bm{f}_{1}^{\top},\cdots,\bm{\tau}_{n}^{\top},\bm{f}_{n}^{\top}]^{\top}\in\mathbb{R}^{n_{u}}$ in the above state-space model. After rearranging (2), (8), and (10), the state space model is written as follows:
@@ III-B Centroidal Composite Inertia Network
For efficient training, we collect data offline from a repetition of two main motions: forward steps and in-place turning, both parameterized accordingly. First, we carefully choose the upper and lower bounds of the feet and base in SE(3) for each motion, as well as the range of swing heights. Then, we uniformly sample initial and final points in SE(3) for each frame at each step and interpolate them to create motion trajectories. For each interpolated data point (input to the network), we solve the inverse kinematics (IK) to compute the joint configurations and centroidal inertia tensors (output to the network). To collect more reliable data, we ensure that no self-collision occur at any interpolated data point. This is done using Pinocchio’s HPP-FCL library on our simplified collision-model of DRACO $3$ involving cubes, spheres, and capsule primitives, for which a self-collision matrix was also configured. Moreover, we also verify that a valid IK solution exists at all instances given the rolling contact joint constraint of the legs of DRACO $3$.
@@ IV-A Experimental Setup
We performed simulations on the DRACO 3 humanoid robot [1] in PyBullet [22]. All control modules employed in this paper are written in C++. The WBC control commands are executed at a frequency of 800 $\mathrm{H}\mathrm{z}$, while the MPC controller operates at 200 $\mathrm{H}\mathrm{z}$. For the execution of contact wrench commands from the MPC, we utilize a Zero-Order Hold (ZOH) between control signals. The MPC prediction horizon consists of 10 knot points, corresponding to 0.5 $\mathrm{s}$.
The centroidal composite inertia network is composed of a 3-layer neural network consisting of a 12-dimensional input layer and two hidden layers (each with 64 neurons and a $\tanh$ activation function). Additionally, DRACO 3’s WBO function approximator is constructed using monomial basis functions up to the 2nd order, with coefficients optimized using the algorithm described in [13] through IPOPT [23]. We simplified the configuration space to 16 dimensions by excluding the ankle, wrist, and neck joints from the optimization, due to their minimal contribution to the CAM. The neural network and WBO function approximator were initially developed using the CasADi symbolic framework [24] in Python and subsequently exported to C++ using CasADi’s internal code generation tool.

### v2

@@ Abstract
This paper proposes a novel control framework for agile and robust bipedal locomotion, addressing model discrepancies between full-body and reduced-order models. Specifically, assumptions such as constant centroidal inertia have introduced significant challenges and limitations in locomotion tasks. To enhance the agility and versatility of full-body humanoid robots, we formalize a Model Predictive Control (MPC) problem that accounts for the variable centroidal inertia of humanoid robots within a convex optimization framework, ensuring computational efficiency for real-time operations. In the proposed formulation, we incorporate a centroidal inertia network designed to predict the variable centroidal inertia over the MPC horizon, taking into account the swing foot trajectories—an aspect often overlooked in ROM-based MPC frameworks. By integrating the MPC-based contact wrench planning with our low-level whole-body controller, we significantly improve the locomotion performance, achieving stable walking at higher velocities that are not attainable with the baseline method. The effectiveness of our proposed framework is validated through high-fidelity simulations using our full-body bipedal humanoid robot, DRACO 3 [1], demonstrating dynamic behaviors.
@@ I INTRODUCTION
Achieving high-speed, agile locomotion is essential for humanoid robots to execute human-level tasks in dynamic environments. Extensive research has been dedicated to enchancing the stability and robustness of bipedal robots during high-speed motions and stationary phases by optimizing balanced body postures [2, 3, 4, 5, 6]. Given the high-dimensional and nonlinear nature of humanoid robots, many of these studies have simplified the locomotion problem using reduced-order models (RoMs) such as the linear inverted pendulum (LIP) [2] or single rigid-body (SRB) models [7]. While these RoMs provide accessible solutions to the locomotion challenges, they often fail to capture critical aspects of robot dynamics, particularly behaviors involving angular momentum. In this paper, we introduce a novel model predictive control (MPC) framework, dubbed Variable Inertia MPC (VI-MPC), which incorporates variable inertia, dependent on the robot’s posture, as opposed to the constant inertia typically employed in conventional methods, to enable more stable and agile locomotion.
I-A Related Works
Model predictive control (MPC) assuming the SRB model has been one of the dominant control strategies applied in quadrupedal locomotion [7, 8, 9, 10]. More recently, this approach has been adapted in humanoid robots [11, 6, 12, 13, 14]. A commonality across these applications is their focus in optimizing locomotion variables such as reaction forces or footsteps under the assumption that the robot can be modeled as a single lumped mass—often informally referred to as the “potato model.” However, in stark contrast to most quadruped robots, this simplification falls short for humanoid robots that feature heavy legs. Specifically, significant changes in the rotational inertia tensor, e.g., resulting from leg swings, compromise the model’s accuracy. Since the MPC’s performance is highly dependent on the precision of the underlying model, the SRB model-based MPC controller becomes less suitable for humanoid robots undergoing high inertia variations.
@@ I-A Related Works
Alternatively, recent works have explored the use of online nonlinear MPC for generating whole-body motions that account for significant changes in angular momentum by employing more expressive models. For example, [17] utilizes the centroidal dynamics and whole-body kinematics optimizers to compute the centroidal and joint trajectories. Similarly, [18] employs a full-body dynamics model to explicitly consider the nonholonomic nature of angular momentum. While these approaches have demonstrated effectiveness, they still suffer from slow computation times and susceptibility to local minima.
I-B Contributions
This paper proposes the VI-MPC framework that accounts for variations in centroidal rotational inertia resulting from leg swing motions. Building on our previous work in [16], we utilize a centroidal composite inertia neural network (CCINN) offline to obtain variable inertial properties. These properties are then seamlessly integrated into our convex MPC formulation to plan locomotion trajectories, leveraging both whole-body orientation (WBO) and the SRB model. The generated trajectories are subsequently tracked using an optimization-based whole-body controller (WBC).
The main contributions of this paper are threefold. First, by incorporating variable inertial properties and WBO into the proposed framework, our approach captures more precise angular dynamics, leading to improved accuracy over conventional methods that rely solely on base link orientation [11, 6, 13]. Second, the combination of offline and online processes in the proposed framework ensures more stable and efficient computations, minimizing the risk of converging to local minima during the optimization process. Lastly, we demonstrate that the proposed MPC approach facilitates more dynamic locomotion, including longer strides, higher walking speeds, and more agile turning behaviors, by effectively bridging the gap between the full-body and SRB models.
The organization of the remainder of this paper is as follows: In Section II, we revisit the dynamics models for control and formally define our problem. Section III details our approach to solving this problem through the MPC and WBC framework. Then, we present a set of numerical simulation results on the MPC in Section IV. Lastly, Section V concludes the paper with potential directions for future work.
@@ II PRELIMINARIES
II-A Full-body Dynamic (FBD) Model
The rigid body dynamics equation of a floating base robot with $n$ degrees of freedom (DoF), actuated by $n_{a}$ joints ($n_{a}<n$), is given by:
$\mathbf{M}(\mathbf{q})\dot{\mathbf{v}}+\mathbf{b}(\mathbf{q},\mathbf{v})=\mathbf{S}_{a}^{\top}\bm{\tau}+\mathbf{J}_{c}(\mathbf{q})^{\top}\mathbf{F}_{r},$ (1)
where $\mathbf{q}=[\mathbf{q}_{f}^{\top},\mathbf{q}_{a}^{\top}]^{\top}\in\mathcal{Q}$, denotes the robot’s configuration, $\mathbf{q}_{f}\in SE(3)$ represents the position and orientation of the system’s floating base, $\mathbf{q}_{a}\in\mathbb{R}^{n_{a}}$ is the actuated joints’ configuration, and $\mathbf{v}\in\mathbb{R}^{n_{a}+6}$ the system’s generalized velocity. $\mathbf{M}(\mathbf{q})\in\mathbb{R}_{>0}^{(n_{a}+6)\times(n_{a}+6)}$, and $\mathbf{b}(\mathbf{q},\mathbf{v})\in\mathbb{R}^{n_{a}+6}$ denote the mass/inertia matrix and the sum of the centrifugal/Coriolis and gravitational force, respectively. The selection matrix for the actuated joint torque vector $\bm{\tau}\in\mathbb{R}^{n_{a}}$ is represented by $\mathbf{S}_{a}\in\mathbb{R}^{n_{a}\times(n_{a}+6)}$. In addition, $\mathbf{F}_{r}\in\mathbb{R}^{n_{c}}$ represents the concatenation of the $6$-dimensional contact wrenches, and $\mathbf{J}_{c}(\mathbf{q})\in\mathbb{R}^{n_{c}\times(n_{a}+6)}$ is the corresponding contact Jacobian. For simple notations, we use $\mathbf{M}$, $\mathbf{b}$, and $\mathbf{J_{c}}$ instead of using $\mathbf{M}(\mathbf{q})$, $\mathbf{b}(\mathbf{q},\dot{\mathbf{q}})$, and $\mathbf{J_{c}(\mathbf{q}})$, respectively.
The dynamics equation in (1) can be decomposed into its unactuated and actuated parts as follows [19]:
$\mathbf{M}_{u}\dot{\mathbf{v}}+\mathbf{b}_{u}$ $=\mathbf{J}_{c,u}^{\top}\mathbf{F}_{r},$ (2a)
$\mathbf{M}_{a}\dot{\mathbf{v}}+\mathbf{b}_{a}$ $=\bm{\tau}+\mathbf{J}_{c,a}^{\top}\mathbf{F}_{r}$ (2b)
where the subscript $u$ and $a$ correspond to unactuated and actuated dynamics, respectively. The unactuated dynamics is equivalent to the robot centroidal momentum dynamics as follows:
$\begin{bmatrix}\mathbf{\dot{l}}_{G}\\ \mathbf{\dot{k}}_{G}\end{bmatrix}=\begin{bmatrix}m\mathbf{g}+\sum_{i=1}^{n}\mathbf{f}_{i}\\ \sum_{i=1}^{n}\mathbf{r}_{i}\times\mathbf{f}_{i}+\bm{\tau}_{i}\end{bmatrix}$ (3)
where $\mathbf{l}_{G}$ and $\mathbf{k}_{G}$ denote the linear and angular momentum expressed at the CoM. $m$ represents the lumped mass of the robot, and $\mathbf{g}\in\mathbb{R}^{3}$ is the gravitational acceleration. $\mathbf{f}_{i}\in\mathbb{R}^{3}$ and $\bm{\tau}_{i}\in\mathbb{R}^{3}$ represent the $i^{\textrm{th}}$ ground reaction force and torque, respectively. $\mathbf{r}_{i}\in\mathbb{R}^{3}$ is the $i^{\textrm{th}}$ footstep position relative to CoM. Note that all the quantities are expressed in the inertial frame unless explicitly stated otherwise.
@@ II-B Single Rigid Body Dynamic Model
ROMs offer the advantage of capturing the simplicity of locomotion principles while enabling the stabilization of robotic systems in fast and reactive manners. The SRB model has been among the most widely used to perform legged robot locomotion. The SRB model includes three major assumptions: 1) the leg dynamics have a negligible impact on the centroidal states of the robot, 2) the utilization of a constant rotational inertia matrix, $\mathbf{I}\in\mathbb{R}^{3\times 3}$, calculated based on the robot’s nominal configuration, and 3) the dynamics of the robot’s base orientation fully capture its overall orientation dynamics. Applying these assumptions to the centroidal momentum dynamics in (3), the equations of motion of the SRB model and its rotational motion are expressed as:
@@
where $\mathbf{p}_{c}\in\mathbb{R}^{3}$ denotes the position of the robot’s Center of Mass (CoM). $\mathbf{R}\in SO(3)$ and $\bm{\omega}\in\mathbb{R}^{3}$ denote the rotation matrix, which transforms from the body frame to world coordinates, and the body angular velocity.
@@ II-C Whole-body Orientation
where $\mathbf{I}_{G}$ represent the the centroidal composite rigid body inertia (CCRBI) [22].
@@ III THE PROPOSED APPROACH
This section introduces our proposed variable inertia MPC-based control framework designed to enable robust dynamic locomotion of humanoid robots, as depicted in Fig 2. We begin this section by introducing the simplified model used in our framework (III-A). Nest, we describe the CCIN (III-B) and the process for generating the reference trajectory (III-C). Following that, we explain the proposed convex MPC formulation (III-D) and the WBC desgined to trck the results of the MPC approach (III-E).
@@ III-A Simplified Model
Inspired by the work of quadruped locomotion control [7, 23], we model the robot as an SRB subject to contact patches at each stance foot. However, there are two key differences in our approach: first, we explicitly take the full-body composite inertia into account along with the SRB. Additionally, instead of the robot’s torso orientation, we employ the WBO dynamics for the SRB orientation dynamics with the following assumption:
Assumption 1.
With the above assumption, we rewrite the SRB rotational dynamics equations (5) and (6) in II-B as follows:
@@
where $\mathbf{I}(\mathbf{q})\in\mathbb{R}^{3\times 3}$ denotes the centroidal composite rigid body inertia (CCRBI) [21], which is a function of joint configuration $\mathbf{q}$. $\mathbf{R}_{\textrm{WBO}}$ represents the rotational centroidal orientation of a multibody system, while $\bm{\omega}_{\textrm{WBO}}\in\mathbb{R}^{3}$ denotes the system’s whole body angular velocity. Given $\mathbf{T}(\bm{\Theta})\in\mathbb{R}^{3\times 3}$, which is the matrix that maps from Euler angle rates to angular velocities, we introduce an assumption as follows:
Assumption 2.
In our problem, we assume the matrix $\mathbf{T}$ is always invertible, namely $\phi\neq\frac{\pi}{2}$ (robot’s WBO is not pointed vertically, which in practice does not happen).
Then, for state-space representation, the rotational kinematics in (11) is rewritten using Euler angles as follows:
@@
where $\bm{\Theta}_{\textrm{WBO}}=[\theta,\phi,\psi]$ follows an earth-fixed roll($\theta$), pitch($\phi$), yaw($\psi$) convention.
@@
We define the state variable and control input as follows:
$\begin{split}\bm{x}=&[\bm{\Theta}_{\textrm{WBO}}^{\top},\bm{p}_{c}^{\top},\bm{w}_{\textrm{WBO}}^{\top},\dot{\bm{p}}_{c}^{\top}]^{\top}\in\mathbb{R}^{n_{x}},\\ \bm{u}=&[\bm{\tau}_{1}^{\top},\bm{f}_{1}^{\top},\cdots,\bm{\tau}_{n}^{\top},\bm{f}_{n}^{\top}]^{\top}\in\mathbb{R}^{n_{u}}.\end{split}$ (13)
After rearranging (4), (10), and (12) usng the defined state and input variables, the state space model is written as follows:
@@ III-B Centroidal Composite Inertia Network
For efficient training, we collect data offline from a repetition of two main motions: forward steps and in-place turning, both parameterized accordingly. First, we carefully choose the upper and lower bounds of the feet and base in SE(3) for each motion, as well as the range of swing heights. Then, we uniformly sample initial and final points in SE(3) for each frame at each step and interpolate them to create motion trajectories. For each interpolated data point (input to the network), we solve the inverse kinematics (IK) to compute the joint configurations and centroidal inertia tensors (output to the network).
@@ IV-A Experimental Setup
We performed simulations on the DRACO 3 humanoid robot [1] in PyBullet [30]. The robot weighs $39~{}$\mathrm{k}\mathrm{g}$$ and features rolling contact joints in its lower body, enabling a broad spectrum of movements. With a total of $n_{a}=25$ DoF, DRACO $3$ includes $6$ DoF per leg, $6$ DoF per arm, and $1$ DoF in the neck, as shown in [1]. All Experiments were performed on a desktop PC with $48$GB of memory and an Intel® CoreTM i7-7700K 4-core CPU @ 4.20GHz. All control modules employed in this paper are written in C++. The WBC control commands are executed at a frequency of 800 $\mathrm{H}\mathrm{z}$, while the MPC controller operates at 200 $\mathrm{H}\mathrm{z}$. For the execution of contact wrench commands from the MPC, we utilize a Zero-Order Hold (ZOH) between control signals. The MPC prediction horizon consists of 10 knot points, corresponding to 0.5 $\mathrm{s}$.
The centroidal composite inertia network is composed of a 3-layer neural network consisting of a 12-dimensional input layer and two hidden layers (each with 64 neurons and a $\tanh$ activation function). To collect more reliable data for the NN training, we ensure that no self-collision occurs at any interpolated data point. This is done using Pinocchio’s HPP-FCL library on our simplified collision model of DRACO $3$ involving cubes, spheres, and capsule primitives, for which a self-collision matrix was also configured. Moreover, we also verify that a valid IK solution exists at all instances given the rolling contact joint constraint of the legs of DRACO $3$.
DRACO 3’s WBO function approximator is constructed using monomial basis functions up to the 2nd order, with coefficients optimized using the algorithm described in [20] through IPOPT [31]. We simplified the configuration space to 16 dimensions by excluding the ankle, wrist, and neck joints from the optimization, due to their minimal contribution to the CAM. The neural network and WBO function approximator were initially developed using the CasADi symbolic framework [32] in Python and subsequently exported to C++ using CasADi’s internal code generation tool.